ganeshie8 (ganeshie8):

\(11_{11}^2 = ?_{11}\)

mathslover (mathslover):

What is this? @ganeshie8 can u explain this to me?

OpenStudy (anonymous):

\[131_{11}\]

ganeshie8 (ganeshie8):

that looks correct... actually i thought of asking as we move left in base 11, 11^2 becomes 100 ? right ?

OpenStudy (unklerhaukus):

\[121_{11}\]

ganeshie8 (ganeshie8):

@mathslover subscript means its base 11

OpenStudy (unklerhaukus):

@Zekarias check that

OpenStudy (anonymous):

You made algebra mistake I think @UnkleRhaukus

OpenStudy (unklerhaukus):

121+22+1=144

OpenStudy (anonymous):

IS A_11=10_10 ??

OpenStudy (unklerhaukus):

yes @sauravshakya

OpenStudy (anonymous):

Well then the answer must be less than 121

ganeshie8 (ganeshie8):

im getting 121 too.. but the q i thought of asking was different

OpenStudy (unklerhaukus):

\[11^2_{11}=(11+1)^2=12^2=144\] \[144=121+22+1=121_{11}\]

OpenStudy (anonymous):



OpenStudy (anonymous):